# Leveraging the true depth of LLMs

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities at the cost of high compute requirements. Recent studies have demonstrated that intermediate layers in LLMs can be removed or reordered without substantial accuracy loss; however, this insight has not yet been exploited to improve inference efficiency. Leveraging observed layer independence, we propose a novel method that groups consecutive layers into pairs evaluated in parallel, effectively restructuring the computational graph to enhance parallelism. Without requiring retraining or fine-tuning, this approach achieves an inference throughput improvement of 1.05x–1.20x on standard benchmarks, retaining 95%–99% of the original model accuracy. Empirical results demonstrate the practicality of this method in significantly reducing inference cost for large-scale LLM deployment. Additionally, we demonstrate that modest performance degradation can be substantially mitigated through lightweight fine-tuning, further enhancing the method’s applicability.
@@ 1 Introduction
Figure 1: The effect of LP on execution time (4K tokens) and perplexity (measured against RedPajama[1]). The bigger Llama model achieves both better perplexity and faster execution speeds than the smaller Llama model.
The rapid advancement of LLMs has revolutionized Artificial Intelligence applications across industries. However, the ever-increasing computational demands of these models, with parameters often numbering hundreds of billions, present significant commercial challenges. Efficient inference is crucial for organizations that deploy these models at scale, as it directly impacts operational costs, user experience, and environmental sustainability [2, 3, 4]. Monthly cloud computing expenses for LLM inference can reach millions of dollars for high-traffic applications, making optimization techniques essential. In addition, reducing inference latency is critical for real-time applications and for deploying models on devices with more limited compute. Thus, the development and implementation of efficient inference methods has become a key differentiator in the competitive AI market, driving both innovation and profitability.
@@
Figure 2: Comparing a normal transformer block with our layer parallel implementation. In (a) we show a normal Transformer Layer. In (b) we illustrate our layer parallelization, where each column runs separately on one or multiple GPUs. The residuals are gathered and synchronized through a reduce operation across all GPUs. The LayerNorms’ weights are copied from the original blocks, and each divergent path uses the normalization from its original layer.
LLMs have evolved to incorporate architectures with hundreds of layers [5, 6]. These models are constructed from stacked transformer blocks, each comprising attention and feedforward subblocks, with a residual stream traversing the entire architecture to facilitate efficient gradient backpropagation during training. This architectural choice parallels the design principles of ResNets [7], where research has shown that network depth may be partially redundant, allowing layer reordering without significant performance degradation [8]. Recent investigations have revealed similar flexibility in transformer architectures [9], where interventions such as layer removal and swapping are applied without large performance degradations. Although these findings challenge our understanding of LLMs’ true effective depth, their potential for optimizing inference efficiency remains unexplored.
Stemming from this line of work, we set to find approaches that reduce the depth of LLMs. We do so by exploring several interventions to the computational graph of pre-trained LLMs such as shuffling, pruning or merging layers. As a result of this search, we find that multiple consecutive block pairs can be processed in parallel while maintaining acceptable performance across perplexity and In-Context Learning (ICL) benchmarks. We introduce a novel parallelism approach, Layer Parallelism (LP), that enhances inference speed when making use of tensor parallelism with a minimal drop in performance. Furthermore, we show that the performance degradation observed across various benchmarks can be partially mitigated through targeted fine-tuning procedures.
@@
  • We explore the space of interventions on a pre-trained LLM layers, and find that some transformations, such as contiguous parallelization, preserve model performance
  • We find that we can define a parallelization transform on the computational graph of two sequential Transformer layers, and stack this parallelization operation to several sequential pairs of layers without loosing significant ICL performance. Our approach, which we call Layer Parallelism (LP), can be applied to existing Transformer models and does not require retraining.
@@
As a result, the rearrangement of the computational graph allows for faster inference speeds when serving the LLM using tensor parallelism. The speed-up mainly arises from the decrease in the number of all-reduce operations between GPUs. This is because LP uses two all-reduce operations per pair of transformer decoder layers, while normal layers would require four all-reduce operations.
@@ 2 Related work
The effective depth of Deep Networks. Theoretically, given enough width, any feed-forward network with at least one hidden layer can model any function [10]. In practice, it is easier to achieve high expressivity by increasing the model’s depth. However, naively increasing network depth can complicate optimization, since the gradients now have to flow through many layers. To alleviate this problem, ResNets [7] introduced skip connections at regular intervals to allow an easy flow of the gradient to the first layers. Alternatively, Inception [11] explored approaches to boost computational power by adding additional processing units along different parallel pathways in the computational network, rather than just along a single sequential path. A unification of both methods can be found in the Highway Networks [12], where the skip connection of the residual blocks consists of another block of compute. Nowadays, residual connections are ubiquitous in large models.
@@
Figure 3: Changes in perplexity when applying transformations on contiguous stretches of layers. Each one of the five heatmaps above correspond to a transformation of a group of consecutive layer, where the row index $s$ corresponds to the first layer of the group, and the column index $e$ to the last. The color coding indicates how the perplexity—estimated on a subset of RedPajama [1]—is impacted by the corresponding modification of the model. The perplexity for the base Llama 2 7B model is $6.2$. In (a), we shuffle—for each forward—the layers from $s$ to $e$. We can see that many consecutive layers can be shuffled with little impact on the overall perplexity. For instance, shuffling layers $15$ to 2525252510101010 layers in total—raises the perplexity only to $9.1$. In (b), we prune contiguous stretches of layers. We can see that not many blocks can be removed without starting to significantly degrade the perplexity. In (c) we merge contiguous layers. The results with merging are near identical to those for pruning. This reveals there is no advantage in merging layers, most likely a results of averaging matrices not originating from the same initial values. In (d) we run contiguous blocks in parallel. Given the success of shuffling, it makes sense that this approach works well. Running blocks $17$ to $27$ raises the perplexity to $9.3$. Finally, in (e) we run pairs of consecutive layers in parallel. As a result, we can parallelize much longer stretches of layers. For instance, we can apply this transformation from layer $4$ to $29$ and only increase the perplexity to $9.1$. This reduces the depth of the model from $32$ to $19$. This result makes it possible for us to leverage this parallelism for faster inference as we discuss in § 4.
Efficient inference of LLMs. Several complementary approaches exist for enhancing the computational efficiency of large-scale models, primarily through pruning/sparsity, quantization, and parallelism. Pruning[13, 14, 15, 16] constitutes a dimensional reduction methodology that systematically eliminates redundant parameters while preserving model performance, thereby introducing architectural sparsity. This methodology is founded on empirical evidence demonstrating that neural networks frequently exhibit overparameterization, containing numerous weights with negligible contributions to the output. Through sophisticated pruning strategies, the inherent sparsity support in contemporary accelerators can be leveraged to enhance both memory utilization and computational efficiency [17, 18]. Early-exit[19, 20] can be seen as a way of runtime layer-wise pruning, which halts the LLM forward pass when the next token certainty is high in the intermediate layers. This approach can be also thought as a way of reducing the effective depth of the model at test time. In contrast, quantization encompasses the transformation of floating-point numerical representations (predominantly FP32) into reduced-precision integer formats, such as INT8 or INT4 [21, 15, 22]. When implemented on hardware accelerators, these lower-precision representations facilitate superior memory bandwidth utilization, addressing a primary bottleneck in modern large-scale models [23]; moreover, integer-based computations yield enhanced processing speed and substantially improved energy efficiency [24]. Finally, parallelization techniques during inference, such as tensor and pipeline parallelism, enable the distribution of computational workload across multiple accelerators, thereby reducing latency and increasing throughput, although this often requires careful consideration of communication overhead and load balancing [25, 26].
Parallel attention-feedforward fusion. GPT-J[27] introduced a parallel formulation of the transformer decoder layer, executing attention and feedforward sub-blocks concurrently:
$y=x+\text{MHA}(\text{LayerNorm}(x))+\text{MLP}(\text{LayerNorm}(x))$
For models trained with tensor parallelism, this modification halves the number of required all-reduce operations. Additionally, it reduces memory bandwidth usage by eliminating one read and write operation of hidden states from HBM. The input projections for the attention and MLP sub-blocks can also be fused into a single kernel, further increasing arithmetic density. Consequently, training time is reduced by approximately 15% without observable performance degradation. PaLM[28] also adopted this formulation, noting that negative effects from deviating from the standard transformer self-attention diminish with increasing model size. This parallel formulation continues to be employed in more recent LLMs, including Gemini 1.5 Flash[29].
We differ from these approaches by focusing specifically on already-trained models with the standard transformer formulation[6, 30, 31], rather than training a model from scratch. By rearranging the computational graph, we merge two consecutive transformer-decoder layers by first executing both attention subblocks and both feedforward subblocks sequentially (Figure 2). Similar to GPT-J and PaLM, the primary source of acceleration in our method arises from halving the number of all-reduce operations across the layers where we apply LP.
Parallelism via Computational Graph Optimization. Recent research has investigated architectural layer-level optimization strategies to enhance transformer model inference efficiency. The Staircase Transformer [32] implements parallel layer execution with dynamic recurrent computation based on model requirements. Similarly, the Staggering Transformer [33] achieves layer parallelization by connecting layer $l_{k}$ at time step $t$ to both the $(t-1)$ output of layer $l_{k-1}$ and the $t$ output of layer $l_{k-2}$. To the best of our knowledge, no research has addressed the fusion of consecutive layers through tensor parallelism.
3 Effective Depth
[…]
Figure 4: Diagram of transformations applied in § 3. Diagrams (a,b,c,d) represent respectively shuffling, merging, pruning and parallel.
In this section we investigate the effective depth of pretrained LLMs by applying several transformations and measuring the resulting perplexity degradation. We reveal loose dependencies between intermediary layers. The transformations consist of shuffling, merging, and pruning transformer layers. To avoid the combinatorial explosion resulting from considering all the possible subsets of transformer layers, we instead apply our transformations to all the contiguous stretch of layers. If $L=\{\ell_{1},\ldots,\ell_{N}\}$ are the ordered layers, then we apply our transformations to all the sub-list $\{\ell_{i}\}_{i=s}^{e}$ with $1\leq s\leq e\leq N$. Previous works have shown that—at least when considering pruning—the importance of layers is well behaved, with low importance layers close to one another [34], which justifies considering contiguous stretch of layers only.
Shuffling blocks. We start by shuffling contiguous stretches of layers, re-ordering the layers according to random permutations. Results are shown in Fig. 3.(a). While shuffling the early and two last layers significantly raises the perplexity, there are large stretches of blocks which can be shuffled with surprisingly low impact on the perplexity. For instance, one can shuffle the layers $\{\ell_{i}\}_{15\leq i<25}$ of Llama 2 7B and only have a increase in perplexity of $2.9$. This goes against the classical belief that models build deeply hierarchical representations, where features in previous layers are leveraged to build more complex features in later layers. We interpret this as multiple layers working at the same level of abstraction. Using these insights, we define the effective depth of an LLM as the shortest depth required to efficiently leverage existing latent representations without a significant loss of performance. This reveals an important level of layer decoupling within the model.
@@ 3 Effective Depth
$+\text{A}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}})))$
$+\text{F}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$ (1)
$+\text{A}_{k+1}({\bm{x}}+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))))$ (SEQ)
@@
$\hat{{\bm{y}}}={\bm{x}}$ $+\text{A}_{k}({\bm{x}})+\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$ (2)
$+\text{A}_{k+1}({\bm{x}})+\text{F}_{k+1}({\bm{x}}+\text{A}_{k+1}({\bm{x}}))$ (PAR)
@@
Contiguous 2-parallel. Instead of parallelizing long stretches of layers, we experiment with running pairs of consecutive layers in parallel. This springs from the assumption that local ordering matters less than global ordering, i.e. shuffling consecutive layers introduces fewer potential issues than shuffling layers separated by larger distances. As an example, if we apply the proposed transformation to layers $\{\ell_{15},\ell_{16},\ell_{17},\ell_{18},\ell_{19}\}$, it would result in the following process: (1) the two layers $\{\ell_{15},\ell_{16}\}$ process the input in parallel (according to equation (PAR)), (2) the output is forwarded to layers $\{\ell_{17},\ell_{18}\}$ which process it in parallel; finally, in (3) their joint output is fed to layer $\ell_{19}$ which processes it on its own as any normal layer. The effect of such transformation of the compute graph on the perplexity can be seen in Fig. 3(e). Remarkably, it is possible to run wide stretches of consecutive pairs of blocks in parallel with only a minor degradation of perplexity. For instance, one can apply this transformation from layer $4$ to layer $29$ with only a degradation of perplexity of $2.9$, while reducing the model depth from $32$ to $19$. The success of this approach led us to also try running triplets of consecutive layers in parallel, but we found it to perform worse.
Exploring other transformations. We also experiment with pruning (Fig. 3(b)) and merging (Fig. 3(c)). Pruning has already been studied in several prior works [35, 36]. While pruning a layer leads to a larger perplexity increase compared to parallelizing two blocks, it offers the advantage of reducing the model’s parameter count, which results in improved throughput and more efficient memory usage.

### v3

@@ Abstract
The remarkable capabilities of Large Language Models (LLMs) are overshadowed by their immense computational cost. While recent work has shown that many LLM layers can be reordered or even removed with minimal impact on accuracy, these insights have not been translated into significant inference speedups. To bridge this gap, we introduce a novel method that restructures the computational graph by grouping and evaluating consecutive layer pairs in parallel. This approach, requiring no retraining, yields a 1.19x throughput gain on Llama 2 7B while reducing the average benchmark accuracy by only 1.5%. We demonstrate the practical value of this method for large-scale LLM deployment and show that some of the lost accuracy can be recovered with lightweight fine-tuning of the parallelized layers.
@@ 1 Introduction
Figure 1: The effect of LP on execution time (4K tokens) and perplexity (measured against RedPajama (together2023redpajama)).
The rapid advancement of LLMs has revolutionized Artificial Intelligence applications across industries. However, the ever-increasing computational demands of these models, with parameters often numbering hundreds of billions, present significant commercial challenges. Efficient inference is crucial for organizations that deploy these models at scale, as it directly impacts operational costs, user experience, and environmental sustainability (singh2025surveysustainabilitylargelanguage; xu2024surveyresourceefficientllmmultimodal; wu2022sustainableaienvironmentalimplications). Monthly cloud computing expenses for LLM inference can reach millions of dollars for high-traffic applications, making optimization techniques essential. In addition, reducing inference latency is critical for real-time applications and for deploying models on devices with more limited compute.
@@
Figure 2: Comparison of a normal transformer block (a) with our layer parallel implementation (b). Divergent paths in (b) are split across the Tensor Parallel axis (Eq. LP).
Thus, the development and implementation of efficient inference methods has become a key differentiator in the competitive AI market, driving both innovation and profitability.
LLMs have evolved to incorporate architectures with hundreds of layers (gpt4; grattafiori2024llama3herdmodels). These models are constructed from stacked transformer blocks, each comprising attention and feedforward subblocks, with a residual stream traversing the entire architecture to facilitate efficient gradient propagation during training. This architectural choice parallels the design principles of ResNets (he2015deepresiduallearningimage), where research has shown that network depth may be partially redundant, allowing layer reordering without significant performance loss (VeitWB16). Recent investigations have revealed similar flexibility in transformer architectures (lad2024remarkablerobustnessllmsstages), where interventions such as layer removal and swapping are applied without large performance degradations. Although these findings challenge our understanding of LLMs’ true effective depth, their potential for optimizing inference efficiency remains unexplored.
Inspired by this observed layer independence, we investigated several interventions to the computational graph of pre-trained LLMs. Our exploration of layer shuffling, pruning, and merging revealed that multiple consecutive block pairs can be processed in parallel while maintaining accuracy across perplexity and In-Context Learning (ICL) benchmarks. This led us to propose Layer Parallelism (LP), a novel approach that enhances inference speed when performing inference in the Tensor Parallel (TP) regime. LP modifies the computational graph of a pre-trained LLM to reduce the inter-device communication by half, with a minimal drop in model performance. Furthermore, we show that this performance degradation can be partially mitigated through targeted fine-tuning procedures.
@@
  • We explore the space of interventions on the layers of a pre-trained LLM and find that some transformations, such as contiguous parallelization, preserve model performance.
  • We find that we can define a parallelization transform on the computational graph of two sequential Transformer layers, and stack this parallelization operation across several sequential pairs of layers without losing significant ICL performance. Our approach, which we call LP, can be applied to existing Transformer models.
@@ 2 Related work
The effective depth of Deep Networks. Theoretically, given enough width, any feed-forward network with at least one hidden layer can model any function (Pinkus_1999). In practice, it is easier to achieve high expressivity by increasing the model’s depth. However, naively increasing network depth can complicate optimization, since the gradients now have to flow through many layers. To alleviate this problem, ResNets (he2015deepresiduallearningimage) introduced skip connections at regular intervals to allow an easy flow of the gradient to the first layers. Alternatively, Inception (szegedy2014goingdeeperconvolutions) explored approaches to boost computational power by adding additional processing units along different parallel pathways in the computational network, rather than just along a single sequential path. A unification of both methods can be found in the Highway Networks (srivastava2015highwaynetworks), where the skip connection of the residual blocks consists of another block of compute. Nowadays, residual connections are ubiquitous in large models.
Efficient inference of LLMs. Several complementary approaches exist for enhancing the computational efficiency of large-scale models, primarily through pruning/sparsity, quantization, and parallelism. Pruning (LeCun1989OptimalBD; han2015learningweightsconnectionsefficient; han2016deepcompressioncompressingdeep; frantar2023sparsegpt) constitutes a dimensional reduction methodology that systematically eliminates redundant parameters while preserving model performance, thereby introducing architectural sparsity. This methodology is founded on empirical evidence demonstrating that neural networks frequently exhibit overparameterization, containing numerous weights with negligible contributions to the output. Through sophisticated pruning strategies, the inherent sparsity support in contemporary accelerators can be leveraged to enhance both memory utilization and computational efficiency (zhang2020sparch; Spatten21). Early-exit (teerapittayanon2017branchynetfastinferenceearly; zhou2020bertlosespatiencefast) can be seen as a way of runtime layer-wise pruning, which halts the LLM forward pass when the next token certainty is high in the intermediate layers. This approach can also be thought of as a way of reducing the effective depth of the model at test time. In contrast, quantization encompasses the transformation of floating-point numerical representations (predominantly FP32) into reduced-precision integer formats, such as INT8 or INT4 (shen2019qberthessianbasedultra; han2016deepcompressioncompressingdeep; jacob2018quantization). When implemented on hardware accelerators, these lower-precision representations allow for higher FLOPs and better use of the memory bandwidth, addressing a primary bottleneck in modern large-scale models (gholami2024ai); moreover, integer-based computations yield enhanced processing speed and substantially improved energy efficiency (horowitz20141). Finally, parallelization techniques during inference, such as tensor and pipeline parallelism, enable the distribution of computational workload across multiple accelerators, thereby reducing latency and increasing throughput, although this often requires careful consideration of communication overhead and load balancing (li2024tpillmserving70bscalellms; narayanan2021efficientlargescalelanguagemodel).
Tensor-parallel optimizations. Inter-device communication remains the dominant bottleneck in tensor parallelism, as each transformer sub-module imposes a synchronization step. Recent work targets this cost by either removing synchronization points or shrinking the communicated tensors. Sync-Point Drop (kim2025spdsyncpointdropefficient) cuts the number of all-reduce operations by modifying block structure so that local attention outputs propagate without immediate aggregation; layers are ranked by sensitivity to synchronization removal, and only the sensitive ones receive targeted tuning, yielding a 20% speedup with roughly 1% accuracy loss. A complementary direction reduces communication volume. (dong2024lowbitcommunicationtensorparallel) quantizes per-layer activation exchanges, reaching a 3.8× compression ratio with about a 2% drop on the evaluated benchmarks.
Parallel attention-feedforward fusion. GPT-J (gptj) introduced a parallel formulation of the transformer decoder layer, executing attention and feedforward sub-blocks concurrently:
$y$ $=x+\operatorname{MHA}(\operatorname{LN}_{\mathrm{MHA}}(x))+\operatorname{FFN}(\operatorname{LN}_{\mathrm{FFN}}(x))$
For models trained with tensor parallelism, this modification halves the number of required all-reduce operations. Additionally, it reduces memory bandwidth usage by eliminating one read and write operation of hidden states from HBM. The input projections for the attention and MLP sub-blocks can also be fused into a single kernel, further increasing arithmetic density. Consequently, training time is reduced by approximately 15% without observable performance degradation. PaLM (chowdhery2023palm) also adopted this formulation, noting that negative effects from deviating from the standard transformer self-attention diminish with increasing model size. This parallel formulation continues to be employed in more recent LLMs, including Gemini 1.5 Flash (team2024gemini).
In contrast to these methods, which require training a new model from scratch with a modified architecture, our approach is applied post-hoc to already-trained models (touvron2023llama2; grattafiori2024llama3herdmodels; yang2025qwen3technicalreport). This highlights two other key differences. First, the granularity of parallelism differs: GPT-J-style models parallelize the attention and feed-forward sub-blocks within a single layer, whereas our method parallelizes entire consecutive layers, directly reducing the model’s effective depth. Second, LP accepts a trade-off by approximating the original computation, which results in a slight performance degradation in exchange for inference acceleration. This degradation can be largely recovered with light fine-tuning. The GPT-J architecture, by contrast, is exact by definition, as the model was trained with it from the beginning.
Parallelism via Computational Graph Optimization. Recent research has investigated architectural layer-level optimization strategies to enhance transformer model inference efficiency. The Staircase Transformer (cutler2025stagformertimestaggeringtransformer) implements parallel layer execution with dynamic recurrent computation based on model requirements. Similarly, the Staggering Transformer (cai2024medusasimplellminference) achieves layer parallelization by connecting layer $l_{k}$ at time step $t$ to both the (t1)(t-1) output of layer $l_{k-1}$ and the $t$ output of layer $l_{k-2}$. To the best of our knowledge, no research has addressed the fusion of consecutive layers through tensor parallelism.
3 Effective Depth
[…]
Figure 3: Diagram of transformations applied in § 3. Diagrams (a,b,c,d) represent shuffling, merging, pruning and parallel respectively.
In this section we investigate the effective depth of pretrained LLMs by applying several transformations and measuring the resulting perplexity degradation. We reveal loose dependencies between intermediate layers. The transformations consist of shuffling, merging, and pruning transformer layers. To avoid the combinatorial explosion resulting from considering all possible subsets of transformer layers, we instead apply our transformations to all contiguous stretches of layers. If $L=\{\ell_{1},\ldots,\ell_{N}\}$ are the ordered layers, then we apply our transformations to all the sublists $\{\ell_{i}\}_{i=s}^{e}$ with $1\leq s\leq e\leq N$. Previous works have shown that—at least when considering pruning—the importance of layers is well-behaved, with low-importance layers close to one another (men2024shortgptlayerslargelanguage), which justifies considering contiguous stretches of layers only.
Shuffling, pruning and merging blocks. We start by investigating the effect of several transformations on the model’s perplexity. First, we experiment with shuffling contiguous stretches of layers (Fig. 3a), re-ordering them according to random permutations. Results, shown in Fig. 4(a), reveal that while shuffling the early and late layers is detrimental, there are large stretches of intermediate blocks that can be shuffled with surprisingly low impact on perplexity. For instance, one can shuffle layers 15 through 24 of Llama 2 7B and only increase perplexity by 2.9. This suggests that many layers may operate at a similar level of abstraction, challenging the classical belief of strictly hierarchical representations. This observed layer decoupling is a key insight. We also experiment with pruning (Fig. 3c) and merging (Fig. 3b) contiguous layers. Pruning, studied in prior works (gromov2024unreasonableineffectivenessdeeperlayers; jung2019compactassessingcompactnessrepresentations), involves removing layers entirely. Merging involves averaging the weights of consecutive layers. We find that both transformations lead to a more significant perplexity increase compared to shuffling (see Fig. 4b and Fig. 4c). Merging, in particular, offers no advantage over pruning, suggesting that naively combining weights from different layers is ineffective. These initial experiments indicate that while layers are robust to reordering, their individual parameters are crucial.
@@ 3 Effective Depth
Figure 4: Changes in perplexity when applying transformations on contiguous stretches of layers. Each of the five heatmaps above corresponds to a transformation of a group of consecutive layers, where the row index $s$ corresponds to the first layer of the group, and the column index $e$ to the last. The color coding indicates how the perplexity—estimated on a subset of RedPajama (together2023redpajama)—is impacted by the corresponding modification of the model. The perplexity for the base Llama 2 7B model is $6.2$. In (a), we shuffle—for each forward—the layers from $s$ to $e$. We can see that many consecutive layers can be shuffled with little impact on the overall perplexity. For instance, shuffling layers $15$ to 25251010 layers in total—raises the perplexity only to $9.1$. In (b), we prune contiguous stretches of layers. We can see that not many blocks can be removed without starting to significantly degrade the perplexity. In (c) we merge contiguous layers. The results with merging are nearly identical to those for pruning. This reveals there is no advantage in merging layers, most likely a result of averaging matrices that originate from different initial values. In (d) we run contiguous blocks in parallel. Given the success of shuffling, it makes sense that this approach works well. Running blocks $17$ to $27$ raises the perplexity to $9.3$. Finally, in (e) we run pairs of consecutive layers in parallel. As a result, we can parallelize much longer stretches of layers. For instance, we can apply this transformation from layer $4$ to $29$ and only increase the perplexity to $9.1$. This reduces the depth of the model from $32$ to $19$. This result makes it possible for us to leverage this parallelism for faster inference as we discuss in § 4.
@@
$+\text{A}_{k+1}({\bm{x}}+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}})))$
$+\text{F}_{k+1}({\bm{x}}+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))$ (1)
$+\text{A}_{k+1}({\bm{x}}+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{A}_{k}({\bm{x}})+\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))))$ (SEQ)
@@
$\hat{{\bm{y}}}$ $={\bm{x}}+\text{A}_{k}({\bm{x}})+\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}}))+\text{A}_{k+1}({\bm{x}})+\text{F}_{k+1}({\bm{x}}+\text{A}_{k+1}({\bm{x}}))$ (PAR)
$\approx{\bm{x}}+\text{A}_{k}({\bm{x}})+\text{A}_{k+1}({\bm{x}})+\text{F}_{k}({\bm{x}}+\text{A}_{k}({\bm{x}})+\text{A}_{k+1}({\bm{x}}))+\text{F}_{k+1}({\bm{x}}+\text{A}_{k}({\bm{x}})+\text{A}_{k+1}({\bm{x}}))$ (LP)
@@
To assess how strongly attention and FFN sub-blocks rely on the residual stream generated by preceding layers, we apply a CKA-based comparison (kornblith2019similarityneuralnetworkrepresentations). For each prompt drawn from a small RedPajama subset, we record module activations under two conditions: a standard forward pass and a counterfactual pass in which the incoming residual contribution is removed before processing the next block. The standard decoder update for layer $k$ is:
$h_{k}$ $=x+A_{k}(x)+F_{k}\big(x+A_{k}(x)\big)$
$A_{k+1}$ $=A_{k+1}(h_{k})$
$F_{k+1}$ $=F_{k+1}\big(h_{k}+A_{k+1}\big)$
$h_{k+1}$ $=h_{k}+A_{k+1}+F_{k+1}$
Counterfactual activations are obtained by subtracting the residual stream before evaluating the next block:
$\tilde{h}_{k}$ $=h_{k}-A_{k}(x)-F_{k}(x+A_{k}(x))$
$\tilde{A}_{k+1}$ $=A_{k+1}(\tilde{h}_{k})$
$\tilde{F}_{k+1}$ $=F_{k+1}(\tilde{h}_{k}+\tilde{A}_{k+1})$
Similarity between original and counterfactual activations is then computed via CKA:
$\mathrm{S}_{A}^{k}$ $=\operatorname{CKA}(A_{k},\tilde{A}_{k})$
$\mathrm{S}_{F}^{k}$ $=\operatorname{CKA}(F_{k},\tilde{F}_{k})$
As shown in Fig. 5, layers that tolerated the 2-parallel intervention also show high CKA similarity, indicating limited dependence of their attention and FFN computations on the immediate residual input.
Figure 5: CKA similarity for Qwen3-4B between the original MHA/FFN activations and the counterfactual activations that exclude incoming residuals. Higher values imply greater invariance to the upstream residual stream. A plateau of high CKA similarity between pairs of layers is preceded by a sharp similarity decline at layer 16, which coincides with the performance degradations experienced when applying different levels of LP at different positions.
Contiguous 2-parallel. Instead of parallelizing long stretches of layers, we experiment with running pairs of consecutive layers in parallel. This springs from the assumption that local ordering matters less than global ordering, i.e. shuffling consecutive layers introduces fewer potential issues than shuffling layers separated by larger distances. As an example, if we apply the proposed transformation to layers $\{\ell_{15},\ell_{16},\ell_{17},\ell_{18},\ell_{19}\}$, it would result in the following process: (1) the two layers $\{\ell_{15},\ell_{16}\}$ process the input in parallel (according to equation (PAR)), (2) the output is forwarded to layers $\{\ell_{17},\ell_{18}\}$ which process it in parallel; finally, in (3) their joint output is fed to layer $\ell_{19}$ which processes it on its own as any normal layer. The effect of such a transformation on the compute graph can be seen in Fig. 4(e). Remarkably, it is possible to run wide stretches of consecutive pairs of blocks in parallel with only a minor degradation of perplexity. For instance, one can apply this transformation from layer $4$ to layer $29$ with only a perplexity increase of $2.9$, while reducing the model depth from $32$ to $19$. The success of this approach led us to also try running triplets of consecutive layers in parallel, but we found it to perform worse.